OpenStudy (anonymous):

Write an inequality that represents the statement "x is greater than –3 and less than or equal to 4." @Cal3blikespie

OpenStudy (anonymous):

here is a template \[___ < x \le ____\]

OpenStudy (anonymous):

\le is less than or equal to

OpenStudy (anonymous):

that should work for most inequalities

OpenStudy (anonymous):

i need tha answer i dont know this @ms_instrumental

OpenStudy (anonymous):



OpenStudy (anonymous):

so this (<) means greater than, the nuber behind the symbol is less than the one in front of it. this one (>) is less than, it has a bigger number behind it than it does in front.

OpenStudy (anonymous):

if x is greater than a number, you put that one first. so -3 < x. than you place what x is less than, in front. so x \le 4. all together, its -3 < x \le 4

OpenStudy (anonymous):

If the number is also equal to then it will have a line under the number. If something is something is "less than x" the problem will look as shown "x>" if something were greater than "x<" the same rules apply when something is also equal to. so you would add a line under the sign to show that it could also be equal to

OpenStudy (anonymous):

hope this helped